How many pounds of nuts selling for $9 a pound must be mixed with 12 pounds of chocolate selling for $5 a pound to create a mixture that sells for $6 a pound?
Answer by rfer(16322)

9x+5(12+x)=6(12+x)
9x+60+5x=72+6x
8x=12
x=1.5 lbs of $9 nuts